# Towards LifeSpan Cognitive Systems

### v1

@@ 4.1 Saving $\mathcal{E}$ as Organized Text
Chunking $\mathcal{E}$ into pieces. Creating a text-based knowledge base involves chunking $\mathcal{E}$ into smaller, manageable units for retrieval. Common chunking strategies divide the text into fixed-sized segments based on token count [36, 42, 46, 94]. Determining the optimal chunk size is crucial for effective retrieval. Large chunks (e.g., 512 tokens) can capture more context but may include irrelevant information, making it harder to pinpoint specific answers. Small chunks (e.g., 128 tokens) are more concise but may miss important contextual details necessary for accurate retrieval. Depending on the specific dataset requirements, chunking methods range from fine-grained sentence retrieval [16, 17, 50] to coarse-grained document or group of documents retrieval [93, 59, 130, 51].
Updating the text-based knowledge base. In LSCS, the knowledge base is updated as new experiences are encountered. This process can be done in real-time or periodically, ensuring the newest experiences are incorporated. Methods such as asynchronous re-embedding and re-indexing help keep the stored experiences up-to-date [36, 42, 46, 94]. Open-source tools like LangChain and LlamaIndex facilitate real-time updates by continuously integrating new data into the vector database.
Generating responses according to the knowledge base. When a query is presented, the RAG mechanism retrieves relevant text from the knowledge base, which is then used to augment the LLM’s prompt. This process allows the LLM to generate responses that reflect both the query and the relevant stored experiences. Text-based retrieval systems typically use either sparse keyword-based retrievers, such as BM25 [88], or dense embedding-based retrievers, which capture semantic similarities between the query and the stored knowledge [45, 86]. Fine-tuning these retrievers can further improve their performance in domain-specific tasks [93, 135, 131, 20].
@@ 4.2 Storing Knowledge as Knowledge Graphs
The second method for storing $\mathcal{E}$ involves using knowledge graphs, where experiences are represented as structured data (such as tree or graph). This approach allows for a more relational organization of knowledge, capturing not just the content of experiences but also the relationships between different entities and concepts. Knowledge graphs are particularly well-suited for tasks that require reasoning over connections between pieces of information.
@@
Creating the knowledge graph and Retrieving. To create a knowledge graph, we first need to extract entities and relations (i.e. triplets) from $\mathcal{E}$ and then transform them into graph structures, with nodes representing entities and edges representing relationships. The constructed knowledge graphs allow for richer retrieval possibilities compared to simple text-based storage. For instance, a system can retrieve not just isolated facts but also the relational context in which those facts exist, providing deeper insights during retrieval and generation tasks. In recent works, several methods have been proposed to improve the efficiency and accuracy of graph-based retrieval [1, 27]. For instance, RAPTOR [89] and MemWalker [12] utilize tree-based structures to facilitate retrieval by providing contextual information at various levels of abstraction. THEANINE [60] and AriGraph [3] construct memory graphs to organize memories for efficient retrieval. Additionally, KGP [115] proposes building an index across multiple documents using knowledge graphs. IIER [34] constructs a chunk-interaction graph to capture the internal connections between document chunks. Inspired by the hippocampal indexing theory of human long-term memory, HippoRAG [35] employs an LLM to process passages into knowledge graph triples and leverages the Personalized PageRank algorithm to perform context-based retrieval. G-Retriever [39] presents a retrieval approach for general textual graph tasks by formulating subgraph retrieval as a Prize-Collecting Steiner Tree optimization problem.
@@
Updating the knowledge graph. Updating knowledge graphs in LSCS involves dynamically incorporating new nodes and edges as new experiences are acquired. This ensures that the knowledge graph remains up-to-date and continues to reflect the system’s evolving understanding of its environment. Knowledge graphs can also be linked with external data sources, further enriching the knowledge.
@@
Generating with the retrieved sub-graph. To perform generation, Once the relevant sub-graph is retrieved, it is used to augment the LLM’s prompt for generation. The structured nature of knowledge graphs provides additional benefits during this stage, as the system can leverage the explicit relationships between entities to generate more coherent and contextually informed responses. This can be particularly useful in complex reasoning tasks, where understanding the connections between concepts is as important as retrieving the right information.
@@ 5 Methods of Saving $\mathcal{E}$ into Raw Text and Process All
Transformers [105] is the de facto mainstream backbone architecture for most modern LLMs [1, 103, 104, 106]. To augment the self-attention layers that are position-agnostic with position information, traditional absolute positional encodings provide the absolute position information, usually with the help of a sequence of vectors called position embeddings [105, 57, 55]. They, however, inherently have trouble when the model encounters unseen positions in inputs longer than the training length. Inspired by this limitation, relative positional encoding techniques are proposed to depend on the attention logit function only on the relative distances between tokens instead of their absolute positions. Examples include a learned attention logit bias in T5 [85], Transformer-XL [21], Skyformer [13], Sketching [14] and Sandwich [18], a fixed linear attention decay [81], and rotating query and key sequences based on distances such as RoPE [96, 62], CAPE [65] and XPos [100, 23]. As a representative example widely used in multiple LLMs, RoPE [96] rotates the key and query vectors based on positions before computing the inner product. Specifically, each vector $\mathbf{x}$ (either key or query) is split into pairs of elements $\{(x_{0},x_{1}),(x_{2},x_{3}),\cdots\}$, with each pair interpreted as a 2-dimensional vector. Another example method is AliBi [81], which offsets all attention logits between tokens $i,j$ by a linear term $-m(i-j)$ and becomes $\mathbf{q}_{i}^{\top}\mathbf{k}_{j}-m(i-j)$. To this end, the MPT-7B codes implement an offset matrix as an additive term in attention logits. Both papers claimed to generalize to lengths longer than the training length. However, length generalization failures are still widely observed when directly applied to off-the-shelf Transformer-based LLMs [54]. These models also suffer from overwhelming quadratic computational complexity, and the information lost issue [67, 90], which limits their deployment on practical computing devices and efficacy on extreme lengths.
@@
In view of the limitations in the Transformer architecture, some papers revisit the concept of recurrent networks and structured state space models (SSMs) [32] for both length-generalizable and efficient architectures. Recurrent structures usually have a linear computational complexity with respect to the sequence length, a desirable property for handling long contexts. They also maintain a bounded or sometimes fixed information bottleneck, which naturally prevents feature drift when length increases and can alleviate the generation quality degradation issue. RWKV [79] revisits the traditional recurrent neural networks (RNNs) and demonstrates that, contrary to the common belief, a large enough RNN can exhibit impressive performance comparable to many Transformer-based LLMs. Similarly, Retnet [101] proposes retentive networks, and Memba [31] proposes to use SSMs, which combines more complicated techniques such as gating mechanism and prefix sum to improve the expressiveness and efficiency further. However, the recurrent structure also requires a delicate tradeoff between long-term and short-term memory, and their performance on long-context tasks, especially on lifespan data, which requires information retention on the order of years, is still under debate and investigation.
@@ 5.2 Length Extension Methods for Existing LLMs
Various methods have been proposed to extend the context length of existing LLMs, addressing their length limits and tackling the lifespan data better. One line of work focuses on modifying the attention mechanism without changing the model parameters. [37, 123] propose to modify the attention mechanism without changing the model parameters to enable LLMs to handle infinite context lengths. In particular, LM-Infinite [37] extends LLMs to an extreme length of 200M tokens with $O(n)$ complexity and without degradation in perplexity metric while also showing improvements in downstream tasks. [52] proposes to group tokens into blocks and let each block share a relative position before applying the attention mechanism. They also demonstrate improved performance on long-context tasks, even though the grouping operation is not extendable to infinite size, which upper bounds the extension length.
@@
Another line of work focuses on fine-tuning the model on longer texts to adapt to longer contexts. As many absolute or relative position encoding methods are trained on mathematically pre-defined periodic functions, [84, 24, 68, 49] propose to apply LLMs (fine-tuned or not) on decreased period frequencies (which is equivalent to interpolating position indices) to adapt LLMs to longer sequences. This modification changes the computational features of the model, making fine-tuning necessary to adapt the model to the new context length. Some other papers finetune LLMs with designed attention patterns [76, 136] on long contexts, using neural tangent kernel [80], or with low-rank adaptation(LoRA) [15]. [127] proposes a wait-to-attend mechanism to extend length limits for memory-based Transformers. Other ways of key-value cache selection/eviction methods are investigated in [87, 25, 138]. Similarly, [43, 61] tackles long context by learning to prune, select, or summarize contexts dynamically. Alternatively, context compression methods [91] propose to learn to compress long contexts into shorter embedding sequences. Some work proposes alternative position encodings [95, 137, 143] or landmark token embeddings [69] that enable extendable context limits. Compared with the other methods, especially continual learning, this category of methods needs to expand the memory continuously when the system obtains new experiences, while continual learning could absorb these experiences into the model parameters without expanding anything. These approaches, relying on LLMs to aggregate and process contextual information on the fly, might encounter problems in the face of mutually conflicting instructions or tasks like privacy protection and information retrieval [108]. On the other hand, cognitive systems with an awareness of the ideal situation are expected to better maintain and handle the ongoing tasks with complete interactions.
Where to store How Compressed (1-4) E.Write (1-4) E.Read (1-4)
@@ 6 Benefits and Limitations of the Methods in Each Category
In this section, we examine the strengths and weaknesses of the various technologies mentioned above. (The strengths and weaknesses are summarized in Table 1 and Table 2)
@@
Saving $\mathcal{E}$ into Model Parameters. This category of methods embeds knowledge directly into the model’s weights. While existing methods can inject various forms of knowledge into a language model, they may not suffice as standalone solutions for achieving LSCS. This category includes methods related to model editing and continual learning. Current model editing techniques fall short in editing the model with life experiences which could be more complicated than factual statements, and continual learning methods still struggle with catastrophic forgetting. Anothor fundamental discrepancy between LSCS and continual learning is that LSCS requires updates based on small pieces of information, whereas continual learning necessitates constructing datasets for retraining. Despite these limitations, these methods still offer valuable contributions to LSCS, particularly when there is a need to inject substantial knowledge into the model, adjust its behavior, or update the system to reflect shifts in societal values. As the model updates its parameters with new experiences, new memories are conceptually merged with existing ones. This process leads to automatic Abstraction and Experiences Merging, where the raw information is not preserved in full detail, but rather abstracted and integrated into the model. Such methods face challenges with Long-term Retention and Accurate Recalling due to the inherent risk of catastrophic forgetting in continual learning and failure in sequential model editing (Although some works like WISE [110] claim to support numerous sequential edits by using external modules, these should technically fall under the category of “Saving $\mathcal{E}$ into Explicit Memory ”). The efficiency of writing the experiences into the model parameters (E.Write) is low, as it involves training and back-propagation (even in model editing methods such as [72, 73], backpropagation is needed). In contrast, the efficiency of reading the experiences (E.Read) is high, as the generation speed remains unaffected by updates while the effects of the experiences can still be reflected in generation.
@@
Saving $\mathcal{E}$ into Explicit Memory. Memory-based methods show promise as they allow the model to access information from the distant past by storing $\mathcal{E}$ in memory, without significantly increasing inference costs as history lengthens. These methods (1) often involve some form of abstraction, resulting in the inability to predict exact answers when queried on specific questions, as only the abstracted information from the experiences is stored. (2) may be hard to enable long-term retention as there is either explicit forgetting [116, 140] or implicit forgetting [141]. One possible way to mitigate the forgetting issue is to enlarge the memory pool as studied in MemoryLLM [116]. However, the utilization of the memory pool in MemoryLLM may not be optimal, as it can only achieve around 40 steps (approximately 20k input length) of updates without completely forgetting the earlier knowledge. Despite the current limitations of memory-based methods, we believe memory-based methods are still an important part of achieving a powerful LSCS. In LSCS, one important ability is Abstraction and Experiences Merging. In memory-based methods, there are certain forms of compression when writing the experiences into the memory, which aligns with the idea of abstraction (abstracting the incoming knowledge). However, experiences merging is still under-explored in memory-based methods. Though there are some preliminary explorations [40], how to merge the memories, especially in the hidden space is extremely hard and there are barely any works capable of merging memories in a way that similar memories are put together and even lead to new understandings instead of simply concatenating all memories. As for the ability of Long-term Retention and Accurate Recalling, memory-based methods are inherently not able to recall accurate information as the compression and forgetting mechanisms are part of the design of the memory module, which means some detailed information is destined to be missing. However they can still achieve a long-term retention where the major contour of events could be maintained with details forgotten. The efficiency of the Write process is higher than saving $\mathcal{E}$ into model parameters as it usually requires a forward pass to convert the text into some certain space, while the efficiency of the Read process is lower than saving $\mathcal{E}$ into model parameters as the LLM needs to read from the memory module and put this extracted information in the context to process.
@@
Saving $\mathcal{E}$ into Knowledge Bases for Retrieval. storing $\mathcal{E}$ in knowledge bases is essential for achieving LSCS, enabling the system to accumulate vast amounts of knowledge over time. Whether stored as organized text or knowledge graphs, the information can be retrieved efficiently through RAG methods to inform the generation process. While text-based storage excels at handling large volumes of general information, knowledge graphs offer more structured and relational insights. Together, these methods allow LSCS to retrieve and utilize past experiences effectively, adapting to new situations and challenges based on accumulated knowledge. This category includes methods that store experiences as raw text in a knowledge base or create structured representations, like graphs, by extracting triplets from the input data. When extracting triplets, some detailed information that cannot be represented by factual triplets is inevitably lost, and the nodes and edges in the knowledge graph can be updated to resolve conflicts and merge similar relations. However, these abstractions are fairly naive and only simple conflicts and merging can be handled. As for the knowledge base in the form of organized text, the process usually needs to abstract some knowledge from the experiences to create index such as the embedding of summary of the experiences, where the experiences merging could be even harder than using knowledge graph. Thus we say this category of methods involve Partial Abstraction and Experiences Merging. These methods offer strong Long-term Retention and Accurate Recalling, as the knowledge base can store extensive information. The efficiency of the Write process (E.Write) is higher than the above two categories, as it involves processing the input to construct structured representations which may be faster than converting all text into some other spaces. The efficiency of Read process (E.Read) is lower than the above two, as it requires retrieving and integrating relevant information from the knowledge base. In comparison, Saving $\mathcal{E}$ into model parameters does not require retrieval, and saving $\mathcal{E}$ into explicit memory could potentially skip the need of retrieval such as MemoryLLM [116].
@@
Saving $\mathcal{E}$ into Raw Text and Process All While saving $\mathcal{E}$ as the raw text and putting it into context is conceptually straightforward, it faces significant practical challenges. For instance, humans, are estimated to speak an average of 16k words per day [71], totaling hundreds of millions of words over a lifetime [8]. Consequently, current LLMs usually cannot process all past experiences. Even methods that claim to handle infinitely long contexts struggle with effectively recalling important knowledge from the past [31, 99]. Despite extensive efforts to extend these length limits, obstacles such as computational inefficiency and information forgetting remain unresolved [67], preventing these approaches from being a fundamental solution to LSCS. As for the property Abstraction and Experiences Merging mentioned in Section 1, this line of methods does not involve any abstraction or experiences merging. However, as all the past experiences are stored explicitly in the context, this category of methods is capable of achieving Long-term Retention and Accurate Recalling. The Write process (E.Write) is highly efficient, as it merely involves concatenating new experiences to the existing data. However, the Read process (E.Read) is less efficient, as it requires the language model to process all stored experiences, leading to increased computational demands.
7 Our Suggested Paradigm of LSCS

### v2

@@ 4.1 Saving $\mathcal{E}$ as Organized Text
Chunking $\mathcal{E}$ into pieces. Creating a text-based knowledge base involves chunking $\mathcal{E}$ into smaller, manageable units for retrieval. Common chunking strategies divide the text into fixed-sized segments based on token count (Guu et al., 2020; Huang et al., 2023a; Izacard et al., 2023; Siriwardhana et al., 2022). Determining the optimal chunk size is crucial for effective retrieval. Large chunks (e.g., 512 tokens) can capture more context but may include irrelevant information, making it harder to pinpoint specific answers. Small chunks (e.g., 128 tokens) are more concise but may miss important contextual details necessary for accurate retrieval. Depending on the specific dataset requirements, chunking methods range from fine-grained sentence retrieval (Cheng et al., 2023; 2024; Jiang et al., 2023b) to coarse-grained document or group of documents retrieval (Shi et al., 2023; Khattab et al., 2022; Yu et al., 2022; Jiang et al., 2024).
Updating the text-based knowledge base. In LSCS, the knowledge base is updated as new experiences are encountered. This process can be done in real-time or periodically, ensuring the newest experiences are incorporated. Methods such as asynchronous re-embedding and re-indexing help keep the stored experiences up-to-date (Guu et al., 2020; Huang et al., 2023a; Izacard et al., 2023; Siriwardhana et al., 2022). Open-source tools like LangChain and LlamaIndex facilitate real-time updates by continuously integrating new data into the vector database.
Generating responses according to the knowledge base. When a query is presented, the RAG mechanism retrieves relevant text from the knowledge base, which is then used to augment the LLM’s prompt. This process allows the LLM to generate responses that reflect both the query and the relevant stored experiences. Text-based retrieval systems typically use either sparse keyword-based retrievers, such as BM25 (Robertson et al., 2009), or dense embedding-based retrievers, which capture semantic similarities between the query and the stored knowledge (Izacard et al., 2021; Ram et al., 2021). Fine-tuning these retrievers can further improve their performance in domain-specific tasks (Shi et al., 2023; Zhang et al., 2023b; Zan et al., 2022; Dai et al., 2022).
@@ 4.2 Storing Knowledge as Knowledge Graphs
The second method for storing $\mathcal{E}$ is using knowledge graphs, where experiences are represented as structured data (tree or graph), allowing for a more relational organization of knowledge, capturing not just the content of experiences but also the relationships between different entities and concepts. Knowledge graphs are particularly well-suited for tasks that require reasoning over connections between pieces of information.
@@
Creating the knowledge graph and Retrieving. To create a knowledge graph, we first need to extract entities and relations (i.e. triplets) from $\mathcal{E}$ and then transform them into graph structures, with nodes representing entities and edges representing relationships. The constructed knowledge graphs allow for richer retrieval possibilities compared to simple text-based storage. For instance, a system can retrieve not just isolated facts but also the relational context in which those facts exist, providing deeper insights during retrieval and generation tasks. In recent works, several methods have been proposed to improve the efficiency and accuracy of graph-based retrieval (Achiam et al., 2023; Edge et al., 2024). For instance, RAPTOR (Sarthi et al., 2024) and MemWalker (Chen et al., 2023a) utilize tree-based structures to facilitate retrieval by providing contextual information at various levels of abstraction. THEANINE (Kim et al., 2024) and AriGraph (Anokhin et al., 2024) construct memory graphs to organize memories for efficient retrieval. Additionally, KGP (Wang et al., 2023e) propose building an index across multiple documents using knowledge graphs. IIER (Guo et al., 2024) construct a chunk-interaction graph to capture the internal connections between document chunks. Inspired by the hippocampal indexing theory of human long-term memory, HippoRAG (Gutiérrez et al., 2024) employs an LLM to process passages into knowledge graph triples and leverages the Personalized PageRank algorithm to perform context-based retrieval. G-Retriever (He et al., 2024a) presents a retrieval approach for general textual graph tasks by formulating subgraph retrieval as a Prize-Collecting Steiner Tree optimization problem.
@@
Updating the knowledge graph. Updating knowledge graphs in LSCS involves dynamically incorporating new nodes and edges as new experiences are acquired (Wang et al., 2023e; Guo et al., 2024; Gutiérrez et al., 2024). This ensures that the knowledge graph remains up-to-date and reflects the system’s evolving understanding of its environment. Knowledge graphs can also be further linked with external data sources.
@@
Generating with the retrieved sub-graph. To perform generation, Once the relevant sub-graph is retrieved, it is used to augment the LLM’s prompt for generation (Hu et al., 2024; Chen et al., 2024). The structured nature of knowledge graphs provides additional benefits during this stage, as the system can leverage the explicit relationships between entities to generate more coherent and contextually informed responses. This can be particularly useful in complex reasoning tasks, where understanding the connections between concepts is as important as retrieving the correct information.
@@ 5 Methods of Saving $\mathcal{E}$ into Raw Text and Process All
Transformers Vaswani et al. (2017) is the de facto mainstream backbone architecture for most modern LLMs Achiam et al. (2023); Touvron et al. (2023a; b); Wang & Komatsuzaki (2021). To augment the self-attention layers that are position-agnostic with position information, traditional absolute positional encodings provide the absolute position information, usually with the help of a sequence of vectors called position embeddings (Vaswani et al., 2017; Kenton & Toutanova, 2019; Ke et al., 2020). They, however, inherently have trouble when the model encounters unseen positions in inputs longer than the training length. Inspired by this limitation, relative positional encoding techniques are proposed to depend on the attention logit function only on the relative distances between tokens instead of their absolute positions. Examples include a learned attention logit bias in T5 (Raffel et al., 2020), Transformer-XL (Dai et al., 2019), Skyformer (Chen et al., 2021), Sketching (Chen et al., 2022) and Sandwich (Chi et al., 2023), a fixed linear attention decay Press et al. (2021), and rotating query and key sequences based on distances such as RoPE (Su et al., 2021; Li et al., 2023), CAPE (Likhomanenko et al., 2021) and XPos (Sun et al., 2022; Ding et al., 2023). As a representative example widely used in multiple LLMs, RoPE (Su et al., 2021) rotates the key and query vectors based on positions before computing the inner product. Specifically, each vector $\mathbf{x}$ (either key or query) is split into pairs of elements $\{(x_{0},x_{1}),(x_{2},x_{3}),\cdots\}$, with each pair interpreted as a 2-dimensional vector. Another example method is AliBi (Press et al., 2021), which offsets all attention logits between tokens $i,j$ by a linear term $-m(i-j)$ and becomes $\mathbf{q}_{i}^{\top}\mathbf{k}_{j}-m(i-j)$. To this end, the MPT-7B codes implement an offset matrix as an additive term in attention logits. Both papers claimed to generalize to lengths longer than the training length. However, length generalization failures are still widely observed when directly applied to off-the-shelf Transformer-based LLMs (Kaiokendev, 2023). These models also suffer from overwhelming quadratic computational complexity, and the information lost issue (Liu et al., 2024; Shang et al., 2024), which limits their deployment on practical computing devices and efficacy on extreme lengths.
@@
In view of the limitations in the Transformer architecture, some papers revisit the concept of recurrent networks and structured state space models (SSMs) (Gu et al., ) for both length-generalizable and efficient architectures. Recurrent structures usually have a linear computational complexity with respect to the sequence length, a desirable property for handling long contexts. They also maintain a bounded or sometimes fixed information bottleneck, which naturally prevents feature drift when length increases and can alleviate the generation quality degradation issue. RWKV (Peng et al., ) revisit the traditional recurrent neural networks (RNNs) and demonstrates that, contrary to the common belief, a large enough RNN can exhibit impressive performance comparable to many Transformer-based LLMs. Similarly, Retnet (Sun et al., 2023) propose retentive networks, and Memba (Gu & Dao, 2023) propose to use SSMs, which combines more complicated techniques such as gating mechanism and prefix sum to improve the expressiveness and efficiency further. However, the recurrent structure also requires a delicate tradeoff between long-term and short-term memory (Gu & Dao, 2023; Beck et al., 2024), and their performance on long-context tasks, especially on lifespan data, which requires information retention on the order of years, is still under debate and investigation.
@@ 5.2 Length Extension Methods for Existing LLMs
Various methods have been proposed to extend the context length of existing LLMs, addressing their length limits and tackling the lifespan data better. One line of work focuses on modifying the attention mechanism without changing the model parameters. Han et al. (2024); Xiao et al. (2024) propose to modify the attention mechanism without changing the model parameters to enable LLMs to handle infinite context lengths. In particular, LM-Infinite (Han et al., 2024) extends LLMs to an extreme length of 200M tokens with $O(n)$ complexity and without degradation in perplexity metric while also showing improvements in downstream tasks. Jin et al. (2024) propose to group tokens into blocks and let each block share a relative position before applying the attention mechanism. They also demonstrate improved performance on long-context tasks, even though the grouping operation is not extendable to infinite size, which upper bounds the extension length.
@@
Another line of work focuses on fine-tuning the model on longer texts to adapt to longer contexts. As many absolute or relative position encoding methods are trained on mathematically pre-defined periodic functions, Qu (2023); Ding et al. (2024); Liu et al. (2023); Jiang et al. (2023a) propose to apply LLMs (fine-tuned or not) on decreased period frequencies (which is equivalent to interpolating position indices) to adapt LLMs to longer sequences. This modification changes the computational features of the model, making fine-tuning necessary to adapt the model to the new context length. Some other papers finetune LLMs with designed attention patterns (Oren et al., 2024; Zhang et al., 2024c) on long contexts, using neural tangent kernel (Peng et al., 2023), or with low-rank adaptation(LoRA) (Chen et al., 2023b). Yang & Hua (2024) propose a wait-to-attend mechanism to extend length limits for memory-based Transformers. Other ways of key-value cache selection/eviction methods are investigated in Ren & Zhu (2024); Dong et al. (2024); Zhang et al. (2024f). Similarly, Huang et al. (2023b); Lee et al. (2024) tackle long context by learning to prune, select, or summarize contexts dynamically. Alternatively, context compression methods (Shao et al., 2024) propose to learn to compress long contexts into shorter embedding sequences. Some work proposes alternative position encodings (Song et al., 2023; Zhang et al., 2024e; Zhu et al., 2023) or landmark token embeddings (Luo et al., 2024) that enable extendable context limits. Compared with the other methods, especially continual learning, this category of methods needs to expand the memory continuously when the system obtains new experiences, while the former could absorb these experiences into the model parameters without expanding anything. These approaches, relying on LLMs to aggregate and process contextual information on the fly, might encounter problems in the face of mutually conflicting instructions or tasks like privacy protection and information retrieval (Wang et al., 2024a). On the other hand, cognitive systems with an awareness of the ideal situation are expected to better maintain and handle the ongoing tasks with complete interactions.
Where to store How Compressed $\uparrow$ (1-4) E.Write $\uparrow$ (1-4) E.Read $\uparrow$ (1-4)
@@ 6 Benefits and Limitations of the Methods in Each Category
In this section, we examine the strengths and weaknesses of the various technologies mentioned above (The strengths and weaknesses are summarized in Table 1 and Table 2).
@@
Saving $\mathcal{E}$ into Model Parameters. This category includes model editing and continual learning. Current model editing techniques fall short in editing the model with life experiences which could be more complicated than factual statements, and continual learning methods still struggle with catastrophic forgetting. Despite these limitations, these methods still offer valuable contributions to LSCS, particularly when there is a need to inject substantial knowledge into the model, adjust its behavior, or update the system to reflect shifts in societal values. As the model updates its parameters with new experiences, new memories are conceptually merged with existing ones. This process leads to automatic Abstraction and Experiences Merging, where the raw information is not preserved in full detail, but rather abstracted and integrated into the model. Such methods face challenges with Long-Term Retention and Accurate Recalling due to the inherent risk of catastrophic forgetting in continual learning and failure in sequential model editing (Although some works like WISE (Wang et al., 2024c) claim to support numerous sequential edits by using external modules, these should technically fall under the category of “Saving $\mathcal{E}$ into Explicit Memory ”). The efficiency of writing the experiences into the model parameters (E.Write) is low, as it involves training and back-propagation (even in model editing methods such as Meng et al. (2022; 2023), backpropagation is needed). In contrast, the efficiency of reading the experiences (E.Read) is high, as the generation speed remains unaffected by updates while the effects of the experiences can still be reflected in generation.
@@
Saving $\mathcal{E}$ into Explicit Memory. Memory-based methods show promise as they allow the model to access information from the distant past by storing $\mathcal{E}$ in memory, without significantly increasing inference costs as history lengthens. These methods (1) often involve some form of abstraction, resulting in the inability to predict exact answers when queried on specific questions, as only the abstracted information from the experiences is stored. (2) may be hard to enable long-term retention as there is either explicit forgetting (Wang et al., 2024f; Zhong et al., 2023) or implicit forgetting (Zhou et al., 2023). One possible way to mitigate the forgetting issue is to enlarge the memory pool as studied in MemoryLLM (Wang et al., 2024f). However, the utilization of the memory pool in MemoryLLM may not be optimal, as it can only achieve around 40 steps (approximately 20k input length) of updates without completely forgetting the earlier knowledge. Despite the current limitations of memory-based methods, we believe memory-based methods are still an important part of achieving a powerful LSCS. In LSCS, one important ability is Abstraction and Experiences Merging. In memory-based methods, there are certain forms of compression when writing the experiences into the memory, which aligns with the idea of abstraction (abstracting the incoming knowledge). However, experiences merging is still under-explored in memory-based methods. Though there are some preliminary explorations (He et al., 2024b), how to merge the memories, especially in the hidden space is extremely hard and there are barely any works capable of merging memories in a way that similar memories are put together and even lead to new understandings instead of simply concatenating all memories. As for the ability of Long-Term Retention and Accurate Recalling, memory-based methods are inherently not able to recall accurate information as the compression and forgetting mechanisms are part of the design of the memory module, which means some detailed information is destined to be missing. However they can still achieve a long-term retention where the major contour of events could be maintained with details forgotten. The efficiency of the Write process is higher than saving $\mathcal{E}$ into model parameters as it usually requires a forward pass to convert the text into some certain space, while the efficiency of the Read process is lower than saving $\mathcal{E}$ into model parameters as the LLM needs to read from the memory module and put this extracted information in the context to process.
@@
Saving $\mathcal{E}$ into Knowledge Bases for Retrieval. storing $\mathcal{E}$ in knowledge bases is essential for achieving LSCS, enabling the system to accumulate vast amounts of knowledge over time. The vast amounts of knowledge in the past can be stored as organized text or knowledge graphs, where RAG methods can be used to retrieve the relevant knowledge from the organized text or knowledge graphs (Gutiérrez et al., 2024; Sarthi et al., 2024) to augment the generation process. While text-based storage excels at handling large volumes of general information, knowledge graphs offer more structured and relational insights. Together, these methods allow LSCS to retrieve and utilize past experiences effectively, adapting to new situations and challenges based on accumulated knowledge. This category includes methods that store experiences as raw text in a knowledge base or create structured representations, like graphs, by extracting triplets from the input data. When extracting triplets, some detailed information that cannot be represented by factual triplets is inevitably lost, and the nodes and edges in the knowledge graph can be updated to resolve conflicts and merge similar relations. However, these abstractions are fairly naive and only simple conflicts and merging can be handled. As for the knowledge base in the form of organized text, the process usually needs to abstract some knowledge from the experiences to create index such as the embedding of summary of the experiences, where the experiences merging could be even harder than using knowledge graph. Thus we say this category of methods involve Partial Abstraction and Experiences Merging. These methods offer strong Long-Term Retention and Accurate Recalling, as the knowledge base can store extensive information. The efficiency of the Write process (E.Write) is higher than the above two categories, as it involves processing the input to construct structured representations which may be faster than converting all text into some other spaces. The efficiency of Read process (E.Read) is lower than the above two, as it requires retrieving and integrating relevant information from the knowledge base. In comparison, Saving $\mathcal{E}$ into model parameters does not require retrieval, and saving $\mathcal{E}$ into explicit memory could potentially skip the need of retrieval such as MemoryLLM (Wang et al., 2024f).
@@
Saving $\mathcal{E}$ into Raw Text and Process All While saving $\mathcal{E}$ as the raw text and putting it into context is conceptually straightforward, it faces significant practical challenges. For instance, humans, are estimated to speak an average of 16k words per day (Mehl et al., 2007), totaling hundreds of millions of words over a lifetime (Brandreth, 1980). Consequently, current LLMs usually cannot process all past experiences. Even methods that claim to handle infinitely long contexts struggle with effectively recalling important knowledge from the past (Gu & Dao, 2023; Sun et al., 2024). Despite extensive efforts to extend these length limits, obstacles such as computational inefficiency and information forgetting remain unresolved Liu et al. (2024), preventing these approaches from being a fundamental solution to LSCS. As for the property Abstraction and Experiences Merging mentioned in Section 1, this line of methods does not involve any abstraction or experiences merging. However, as all the past experiences are stored explicitly in the context, this category of methods is capable of achieving Long-Term Retention and Accurate Recalling. The Write process (E.Write) is highly efficient, as it merely involves concatenating new experiences to the existing data. However, the Read process (E.Read) is less efficient, as it requires the language model to process all stored experiences, leading to increased computational demands.
7 Our Proposed Instantiation of LSCS